OpenStudy (anonymous):

-2(csc2θ - cot2θ) - 1 simplify

OpenStudy (larseighner):

\(-2(\csc^2\theta - \cot^2\theta) - 1 \) or \(-2(\csc(2\theta) - \cot(2\theta)) - 1\) ?

OpenStudy (anonymous):

first one

OpenStudy (larseighner):

You can simply with a Pythagorean identity. Do you know them?

OpenStudy (vishweshshrimali5):

You may have a look at this identity: \[\large{\boxed{\csc^2{\theta} = 1+\cot^2{\theta}}}\] Or you may proceed s @LarsEighner is advising you to.

OpenStudy (larseighner):

If you do not know a Pythagorean identity to apply, you could express \(\cos^2\theta\) in other terms. Likewise \(\cot^2\theta\)

OpenStudy (larseighner):

The identity @vishweshshrimali5 gave you is one of three Pythagorean identities that are often used when you have functions "squared"

OpenStudy (anonymous):

okay got it thanks!